A company is analyzing two mutually exclusive projects, S and L, whose cash flows are shown...

  1. A company is analyzing two mutually exclusive projects, S and L, whose cash flows are shown below:

    Year Cash Flow E Cash Flow F
    0 -$1,100 -$1,100
    1    1,000           0
    2       350       300
    3         50     1500


             
    The company's cost of capital is 12 percent, and it can get an unlimited amount of capital at that cost. What is the IRR of the better project?

    12.00%

    15.53%

    18.62%

    19.08%

    20.46%
Answer #1

E:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=1000/1.12+350/1.12^2+50/1.12^3

=$1207.46

NPV=Present value of inflows-Present value of outflows

=$1207.46-$1100

=$107.46(Approx).

F:

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=300/1.12^2+1500/1.12^3

=$1306.83

NPV=Present value of inflows-Present value of outflows

=$1306.83-$1100

=$206.83(Approx).

Hence F is better having higher NPV.

Let irr be x%
At irr,present value of inflows=present value of outflows.

1100 =300/1.0x^2+1500/1.0x^3

Hence x=irr=19.08%(Approx).